3. Z(z) = (1 + + 22); ROC: 121 > 1/2

Solve this using Z transform table and algebraic manipulation of the function. Make it as simple as possible.
Answer #1

Dear students.This Question is about the Signal and System.Solution is provided here.14 APRI Monday 104-261 Wk 16 Answith number 13 the function is as followst ya huu, 222 222 - - n - 1+2+22 - - - 2²+2+2 14 AFriday 108-257 WVk 16 - biroz 99-13 14 +233 - -.-22) Salvat - 3* z=(1-13). 14 APRL 2 Saturday 109-256 Wk 16 2.m) = x(+1-) PuVYLU LY 14 APRU - fz 602 ums ( 204 m. penginNice.Thanks and Regards.